# Inclusive Training Separation and Implicit Knowledge Interaction for Balanced Online Class-Incremental Learning

### v1

@@ 4.2 Implicit Knowledge Interaction
Although our $\mathcal{L}_{DC}$ loss provides a way of knowledge exchange by mixing stream samples and buffer samples when $\alpha\in(0,1)$, relying solely on this mixing strategy cannot resolve the imbalance problem, as in the baseline ER method. Therefore, we propose implicit techniques that enable effective knowledge exchange between the buffer classifier and the stream classifier, namely (i) applying proxy-anchor loss to the training of the stream classifier along with the feature extractor, which supports forward transfer, and (ii) momentum-based updating to the buffer classifier from the stream classifier for backward transfer.
@@
Redesigned Proxy-Anchor Loss.  In our separated learning scheme with DC, the stream classifier may not effectively acquire the knowledge of buffer samples, as it is designed to guarantee such independent learning without any interference. However, the ability of forward transfer is essential in OCIL, where each stream batch contains only a small number of samples. To this end, we suggest using our redesigned version of proxy-anchor loss (PAL) [16] for the stream classifier to implicitly utilize knowledge from buffer features. The main idea is to let each class vector of the stream classifier, denoted as $\mathbf{w}\in\mathbf{W}_{str}$, be a proxy of the class during training with features of each buffer batch $\mathcal{B}_{\mathcal{M}}$ using proxy-anchor loss, replacing randomly initialized proxies in the vanilla proxy-anchor loss. We use $\mathcal{Z}_{\mathcal{M}}$ to denote the set of buffer features corresponding to $\mathcal{B}_{\mathcal{M}}$. For each $\mathcal{Z}_{\mathcal{M}}$, we define our redesigned proxy-anchor loss $\mathcal{L}_{PAL}$ as follows:
@@
$\mathcal{L}_{PAL}=\frac{1}{|\mathbf{W}^{+}_{str}|}\sum_{\mathbf{w}\in\mathbf{W}^{+}_{str}}\log(1+\sum_{\mathbf{z}\in\mathbf{Z}_{\mathcal{M}_{\mathbf{w}}}^{+}}e^{-\gamma(\langle\mathbf{\overline{z}},\mathbf{\overline{w}}\rangle-\delta)})$ (3)
@@
$+\frac{1}{|\mathbf{W}_{str}|}\sum_{\mathbf{w}\in\mathbf{W}_{str}}\log(1+\sum_{\mathbf{z}\in\mathbf{Z}_{\mathcal{M}_{\mathbf{w}}}^{-}}e^{\gamma(\langle\mathbf{\overline{z}},\mathbf{\overline{w}}\rangle+\delta)}),$
@@
where $\delta>0$ is a margin, $\gamma>0$ is a constant scaling factor, determining how strongly pulling or pushing embedding vectors, and $\langle\mathbf{\overline{z}},\mathbf{\overline{w}}\rangle$ represents the cosine similarity between $\mathbf{z}$ and $\mathbf{w}$. $\mathbf{W}^{+}_{str}$ denotes positive proxies corresponding to the classes in the buffer batch. With respect to each proxy $\mathbf{w}$, the set of buffer features $\mathbf{Z}_{\mathcal{M}}$ is also divided into its positive and negative subsets, $\mathbf{Z}_{\mathcal{M}_{\mathbf{w}}}^{+}$ and $\mathbf{Z}_{\mathcal{M}_{\mathbf{w}}}^{-}$, respectively.
@@
Momentum Updating.  Lastly, in order to gradually transfer acquired knowledge from the stream classifier to the buffer classifier, we employ a momentum updating mechanism as:
$\mathbf{W}_{buf}\xleftarrow{}m\mathbf{W}_{buf}+(1-m)\mathbf{W}_{str},$ (4)
where $m$ is a momentum coefficient that controls the rate of update for the buffer classifier, typically set close to $1$ (e.g., 0.99 in our experiments). This update is performed at the end of each training step, following the processing of each mini-batch, which contains a few stream and buffer samples (e.g., 10 of each in our experiments).
4.3 The Overall Process of BOIL
The overall process of BOIL is illustrated in Figure 2. BOIL comprises a shared feature extractor $h$, dual classifiers $f(\mathbf{z};{\mathbf{W}_{str/buf}})$ and implicit knowledge interaction modules. Combining all our components, we present the final loss of our BOIL method as follows:
$\mathcal{L}_{BOIL}=\mathcal{L}_{DC}+\beta\mathcal{L}_{PAL},$ (5)
where $\beta$ is a hyperparameter. Our BOIL method trains a balanced online class-incremental learner through widely used cross-entropy loss together with redesigned proxy-anchor loss. More detailed steps are presented in the Appendix.
5 Experiments
5.1 Experimental Setup
Datasets.  We use three real-world benchmark datasets in image classification. Following [7, 23], Split CIFAR-10 is constructed by splitting CIFAR-10 [26] into 5 disjoint tasks, 2 classes per task. Both Split CIFAR-100 and Split Mini-ImageNet contain 10 disjoint tasks, 10 classes per task, by splitting CIFAR-100 [26] and Mini-ImageNet [27], respectively.
Baselines.  We compare our BOIL method with the following replay-based methods in the OCIL setting: ER [2], GSS [8], MIR [6], A-GEM [28], Gdumb [29], ASER [7], SS-IL [3], ER-DVC [23], ER-ACE [4], OCM [13], PCR [12] and LODE [9]. We also include FINE-TUNE as a non-replay baseline for comparison.
Evaluation Metrics.  For performance assessment, we use three metrics commonly used in OCIL [7, 30, 31]: average accuracy (AA), average forgetting (AF), and average intransigence (AI). Intuitively, the higher AA indicates better overall performance, yet the lower AF and AI imply better stability and better plasticity, respectively. To provide their formal definitions, we first let $a_{k,j}\in[0,1]$ to denote the task-wise classification accuracy for the $j$-th task after learning $k\geq j$ continual tasks. According to the single-head evaluation setup [30], each prediction is made across all classes without being aware of task identification. Then, at the $k$-th task, three metrics are defined as follows [31]: $AA_{k}=\frac{1}{k}\sum_{j=1}^{k}{a_{k,j}},~{}AF_{k}=\frac{1}{k-1}\sum_{j=1}^{k-1}{f_{j,k}},$ and $AI_{k}=\frac{1}{k}\sum_{j=1}^{k}{a_{j}^{*}}-a_{j,j},$ where (i) $f_{j,k}=\operatorname{max}_{i\in\{1,\dots,k-1\}}(a_{i,j}-a_{k,j})$ for $\forall{j<k}$, and (ii) $a_{j}^{*}$ indicates an empirical upper-bound for the task-wise accuracy of the $j$-th task, which is obtained from a purely fine-tuned model with respect to $\mathcal{D}_{j}$ without using any other loss terms. In terms of their value ranges, we have $AA\in[0,1]$ , $AF\in[-1,1]$, and $AI\in[-1,1]$.
Implementation Details.  Following the common architecture in OCIL [11, 23, 12], we utilize Reduced ResNet-18 [32] as the feature extractor. For batch size, each stream batch contains 10 images drawn from the data stream, while 10 samples are randomly retrieved from the buffer to form a buffer batch. In our BOIL method, we set $m=0.99$ in Eq. (4), $\gamma=32$ and $\delta=0.1$ in Eq. (5) for all datasets. For Split CIFAR-100 and Split Mini-ImageNet, $\alpha$ in Eq. (2) and $\beta$ in Eq. (5) are set to $0.5$ and $0.2$, respectively, whereas they are set to 00 and $0.1$ for Split CIFAR-10. We reproduce all the evaluations in a consistent environment, where NVIDIA Geforce 3090 GPU and PyTorch toolbox are utilized. Each measurement in all experimental results is the average along with its corresponding standard deviation over 10 independent runs, where each run shuffles classes when splitting datasets. Full implementation details are presented in the Appendix.
Table 1: Average accuracy at the end of training on three datasets. The best scores are highlighted in boldface, while the runner-up scores are underlined. ‘T’ and ‘F’ indicate two categories: training distinction and feature enhancement, respectively.
@@ 5.1 Experimental Setup
BOIL (ours) 26.3 ±0.8 30.3 ±0.7 33.1 ±1.0 51.5 ±1.0 57.5 ±0.6 61.4 ±1.3 23.9 ±0.6 27.0 ±0.6 28.5 ±0.5
5.2 Performance Comparison
Overall Performance.  Table 1 summarizes overall performance evaluation using three datasets, where we present the average accuracy of each model at the end of training over all tasks with different buffer sizes. Our BOIL consistently achieves the best performance among all compared methods in all settings, mostly followed by PCR [12]. Training distinction methods generally tend to show unsatisfactory performance with clear margins except for LODE [9], compared to BOIL. Either SS-IL [3] or ER-ACE [4] relies on exclusive loss separation, leading to less competitive performance, while LODE manages to further improve the performance by relaxed loss separation. In the category of feature enhancement, PCR takes the best position by leveraging proxies in metric learning, but SCR [11] seems to suffer from too few samples in a mini-batch to train with its supervised contrastive loss. Every method obviously shows better performance as buffer size increases, but our BOIL method takes the best use of knowledge from buffer samples, further enlarging the performance gap with a larger buffer space.
@@ 5.2 Performance Comparison
Figure 4: Balance between stability and plasticity. Each graph plots average intransigence (AI) on the x-axis, which measures plasticity, and average forgetting (AF) on the y-axis, which measures stability. Both AI and AF are equally scaled in $[-1,1]$, with lower values indicating better plasticity and stability, respectively. Different colors represent levels of average accuracy.
@@
Balance of Stability and Plasticity.  To evaluate the balance between plasticity and stability, Figure 4 displays interplay graphs [30], which plot each method based on two key metrics: average forgetting (AF) and average intransigence (AI). We also use different colors to show different levels of average accuracy (AA). Both AF and AI have the same range $[-1,1]$, where lower values represent better stability and plasticity, respectively. Note that although AF and AI can technically be negative, a zero value already indicates no deviation from the ideal performance achievable through joint training with the entire dataset. In all the graphs, our BOIL method is clearly positioned closest to the bottom-left corner, and is represented with the darkest color. This demonstrates that BOIL not only outperforms exiting methods in terms of overall performance, but also consistently achieves the best balance between stability and plasticity. Methods like OCM [13] tend to prioritize stability at the expense of plasticity. In contrast, methods with negative AI values (depicted as circular points), such as vanilla ER [2], excessively focus on plasticity, yet they suffer from severe forgetting with respect to AF, consequently failing to secure high AAs. Overall, maintaining balance is crucial for achieving high performance in OCIL, as more balanced methods tend to have darker-colored points, corresponding to higher AAs, with BOIL leading the way, followed by LODE, PCR and ER-ACE.
5.3 Ablation Studies

### v2

@@ 4.2. Implicit Knowledge Interaction
Although our $\mathcal{L}_{DC}$ loss provides a way of knowledge exchange by mixing stream samples and buffer samples when $\alpha\in(0,1)$, relying solely on this mixing strategy cannot resolve the imbalance problem, as in the baseline ER method. Therefore, we propose implicit techniques that enable effective knowledge exchange between the buffer classifier and the stream classifier, namely (i) redesigning proxy-anchor loss to the training of the stream classifier along with the feature extractor, and (ii) proxy alignment feedback to the buffer classifier.
@@
Buffer-to-Stream: Redesigned Proxy-Anchor Loss. In our separated learning scheme with DC, the stream classifier may not effectively acquire the knowledge of buffer samples, as it is designed to guarantee such independent learning without any interference. However, forward transfer is essential in OCIL, where each stream batch contains only a small number of samples. To this end, we suggest using our redesigned version of proxy-anchor loss (PAL) (PAL/KimKCK20) for the stream classifier to implicitly utilize knowledge from buffer features. The main idea is to let each class vector of the stream classifier, denoted as $\mathbf{w}\in\mathbf{W}_{str}$, be a prototype (i.e., proxy) of the class during training with features of each buffer batch $\mathcal{B}_{\mathcal{M}}$ using proxy-anchor loss, replacing randomly initialized proxies in the vanilla proxy-anchor loss. We use $\mathcal{Z}_{\mathcal{M}}$ to denote the set of buffer features corresponding to $\mathcal{B}_{\mathcal{M}}$. For each $\mathcal{Z}_{\mathcal{M}}$, we define our redesigned proxy-anchor loss $\mathcal{L}_{PAL}$ as follows:
@@
(3) $\mathcal{L}_{PAL}=$ $\frac{1}{|\mathbf{W}_{str}^{+}|}\sum_{\mathbf{w}\in\mathbf{W}_{str}^{+}}\log(1+\sum_{\mathbf{z}\in\mathcal{Z}_{M_{w}}^{+}}e^{-\gamma\left(\text{cos}(\mathbf{z},\mathbf{w})-\delta\right)})$
@@
$+\frac{1}{|\mathbf{W}_{str}|}\sum_{\mathbf{w}\in\mathbf{W}_{str}}\log(1+\sum_{\mathbf{z}\in\mathcal{Z}_{M_{w}}^{-}}e^{\gamma\left(\text{cos}(\mathbf{z},\mathbf{w})+\delta\right)})$
@@
where $\delta>0$ is a margin, $\gamma>0$ is a constant scaling factor, determining how strongly pulling or pushing embedding vectors, and $\text{cos}(\mathbf{z},\mathbf{w})$ represents the cosine similarity between $\mathbf{z}$ and $\mathbf{w}$. $\mathbf{W}^{+}_{str}$ denotes positive proxies corresponding to the classes in the buffer batch. For each proxy $\mathbf{w}$, the set of buffer features $\mathcal{Z}_{\mathcal{M}}$ is also divided into its positive and negative subsets, $\mathcal{Z}_{\mathcal{M}_{\mathbf{w}}}^{+}$ and $\mathcal{Z}_{\mathcal{M}_{\mathbf{w}}}^{-}$, respectively.
@@
Table 1. Average accuracy at the end of training on three datasets. The best scores are highlighted in boldface, while the runner-up scores are underlined. ‘T’ and ‘F’ indicate two categories: training distinction and feature enhancement, respectively.
@@
BISON (ours) 26.3 ±1.0 30.0 ±0.7 32.8 ±0.9 52.5 ±0.7 58.0 ±1.0 61.2 ±0.6 24.2 ±0.4 26.9 ±0.6 28.7 ±0.3
Stream-to-Buffer: Proxy Alignment Feedback. For knowledge transfer from the stream classifier to the buffer classifier, we focus on the fact that the weight vector corresponding to each class in both classifiers technically serves as the class prototype. Based on this insight, we suggest proxy alignment feedback (PAF), that transfers the prototype-level information enhanced by our redesigned proxy-anchor loss. Specifically, given the set of class labels $\mathcal{Y}_{B^{s-1}_{\mathcal{M}}}$ observed in the previous buffer batch, we align the corresponding proxy weights between two classifiers as shown below:
(4) $\mathcal{L}_{Align}=\frac{1}{|\mathcal{Y}_{B^{s-1}_{\mathcal{M}}}|}\sum_{c\in\mathcal{Y}_{B^{s-1}_{\mathcal{M}}}}\Big(1-\cos\!\big(\mathbf{W}^{s}_{buf}[c,:],\;\mathbf{W}^{s-1}_{str}[c,:]\big)\Big),$
where $\text{cos}(\cdot,\cdot)$ symbolizes cosine similarity as mentioned before and $\mathbf{W}^{s-1}_{str}[c,:]$ is frozen. The refined weights in $\mathbf{W}_{str}$ from the previous step are treated as fixed teacher proxies and are aligned with their buffer counterparts $\mathbf{W}_{buf}$. Note that the $\mathcal{L}_{Align}$ is calculated before the current step’s back-propagation, both classifiers are evaluated at the same temporal state, while the gradient is restricted to $\mathbf{W}_{buf}$.
4.3. The Overall Process of BISON
The overall process of BISON is illustrated in Figure 2. BISON comprises a shared feature extractor $h$, dual classifiers $f(\mathbf{z};{\mathbf{W}_{str/buf}})$ and implicit knowledge interaction modules. Combining all our components, we present the final loss of our BISON method as follows:
(5) $\mathcal{L}_{BISON}=\mathcal{L}_{DC}+\beta\mathcal{L}_{PAL}+\lambda\mathcal{L}_{Align},$
where $\beta$ and $\lambda$ are hyper-parameters. BISON trains a balanced online class-incremental learner through widely used cross-entropy loss together with redesigned proxy-anchor loss and proxy alignment feedback. More detailed steps are presented in the Appendix.
5. Experiments
5.1. Experimental Setup
Datasets. We use three real-world benchmark datasets in image classification. Following (ASER/ShimMJSKJ21; DVC/Gu0WD22), Split CIFAR-10 is constructed by splitting CIFAR-10 (CIFAR/krizhevsky2009learning) into 5 disjoint tasks, 2 classes per task. Both Split CIFAR-100 and Split Mini-ImageNet contain 10 disjoint tasks, 10 classes per task, by splitting CIFAR-100 (CIFAR/krizhevsky2009learning) and Mini-ImageNet (Mini/VinyalsBLKW16), respectively.
Baselines. We compare our BISON method with the following replay-based methods in the OCIL setting: ER (ER/chaudhry2019tiny), GSS (GSS/AljundiLGB19), MIR (MIR/AljundiBTCCLP19), A-GEM (AGEM/chaudhry2018efficient), Gdumb (Gdumb/PrabhuTD20), ASER (ASER/ShimMJSKJ21), SS-IL (SSIL/AhnKLBKM21), ER-DVC (DVC/Gu0WD22), ER-ACE (ACE/CacciaAATPB22), OCM (OCM/GuoLZ22), PCR (PCR/LinZFLY23) and LODE (LODE/LiangL23). We also include FINE-TUNE as a non-replay baseline for comparison.
Evaluation Metrics. For performance assessment, we use three metrics commonly used in OCIL (ASER/ShimMJSKJ21; Interplay/ChaudhryDAT18; CPR/ChaHHCM21): average accuracy (AA), average forgetting (AF), and average intransigence (AI). Intuitively, the higher AA indicates better overall performance, yet the lower AF and AI imply better stability and better plasticity, respectively. To provide their formal definitions, we first let $a_{k,j}\in[0,1]$ to denote the task-wise classification accuracy for the $j$-th task after learning $k\geq j$ continual tasks. According to the single-head evaluation setup (Interplay/ChaudhryDAT18), each prediction is made across all classes without being aware of task identification. Then, at the $k$-th task, three metrics are defined as follows (CPR/ChaHHCM21): $AA_{k}=\frac{1}{k}\sum_{j=1}^{k}{a_{k,j}},~AF_{k}=\frac{1}{k-1}\sum_{j=1}^{k-1}{f_{j,k}},$ and $AI_{k}=\frac{1}{k}\sum_{j=1}^{k}{a_{j}^{*}}-a_{j,j},$ where (i) $f_{j,k}=\operatorname{max}_{i\in\{1,\dots,k-1\}}(a_{i,j}-a_{k,j})$ for $\forall{j<k}$, and (ii) $a_{j}^{*}$ indicates an empirical upper-bound for the task-wise accuracy of the $j$-th task, which is obtained from a purely fine-tuned model with respect to $\mathcal{D}_{j}$ without using any other loss terms. In terms of their value ranges, we have $AA\in[0,1]$ , $AF\in[-1,1]$, and $AI\in[-1,1]$.
Implementation Details. Following the common architecture in OCIL (SCR/MaiLKS21; DVC/Gu0WD22; PCR/LinZFLY23), we utilize Reduced ResNet-18 (ResNet/HeZRS16) as the feature extractor. For batch size, each stream batch contains 10 images drawn from the data stream, while 10 samples are randomly retrieved from the buffer to form a buffer batch. In our BISON method, we keep the default setting as (PAL/KimKCK20) does where $\gamma=32$ and $\delta=0.1$ in Eq. (5) for all datasets. For Split CIFAR-100 and Split Mini-ImageNet, $\beta$ and $\lambda$ in Eq. (5) are set to $0.2$ and $10.0$, respectively, whereas they are set to $0.1$ and $3.0$ for Split CIFAR-10. We reproduce all the evaluations in a consistent environment, where NVIDIA Geforce 3090 GPU and PyTorch toolbox are utilized. Each measurement in all experimental results is the average along with its standard deviation over 10 independent runs, where each run shuffles classes when splitting datasets. Full implementation details and hyperparameter determination are presented in the Appendix.
@@ 5.1. Experimental Setup
Figure 4. Balance between stability and plasticity. Each graph plots average intransigence (AI) on the x-axis, which measures plasticity, and average forgetting (AF) on the y-axis, which measures stability. Both AI and AF are equally scaled in [1,1][-1,1], with lower values indicating better plasticity and stability, respectively. Different colors represent levels of average accuracy.
5.2. Performance Comparison
Overall Performance. Table 1 summarizes overall performance evaluation using three datasets, where we present the average accuracy of each model at the end of training over all tasks with different buffer sizes. Our BISON consistently achieves the best performance among all compared methods in all settings, mostly followed by PCR (PCR/LinZFLY23). Training distinction methods generally tend to show unsatisfactory performance with clear margins except for LODE (LODE/LiangL23), compared to BISON. Either SS-IL (SSIL/AhnKLBKM21) or ER-ACE (ACE/CacciaAATPB22) relies on exclusive loss separation, leading to less competitive performance, while LODE manages to further improve the performance by relaxed loss separation. In the category of feature enhancement, PCR takes the best position by leveraging proxies in metric learning, but SCR (SCR/MaiLKS21) and OCM (OCM/GuoLZ22) seem to suffer from too few samples in a mini-batch to train with their contrastive loss. We also examine the case of the larger buffer batch in the Appendix. Every method obviously shows better performance as buffer size increases, but our BISON method takes the best use of knowledge from buffer samples, further enlarging the performance gap with a larger buffer space. We also report the incremental performance over the steps in Figure A1 of Appendix, where our BISON method generally shows the highest performance throughout the entire training process.
@@ 5.2. Performance Comparison
Balance of Stability and Plasticity. To evaluate the balance between plasticity and stability, Figure 4 displays interplay graphs (Interplay/ChaudhryDAT18), which plot each method based on two key metrics: average forgetting (AF) and average intransigence (AI). We also use different colors to show different levels of average accuracy (AA). Both AF and AI have the same range [1,1][-1,1], where lower values represent better stability and plasticity, respectively. Note that although AF and AI can technically be negative, a zero value already indicates no deviation from the ideal performance achievable through joint training with the entire dataset. In all the graphs, our BISON method is clearly positioned closest to the bottom-left corner, and is represented with the darkest color. This demonstrates that BISON not only outperforms exiting methods in terms of overall performance, but also consistently achieves the best balance between stability and plasticity. Methods like OCM (OCM/GuoLZ22) tend to prioritize stability at the expense of plasticity. In contrast, methods with negative AI values (depicted as circular points), such as vanilla ER (ER/chaudhry2019tiny), excessively focus on plasticity, yet they suffer from severe forgetting with respect to AF, consequently failing to secure high AAs. Overall, maintaining balance is crucial for achieving high performance in OCIL, as more balanced methods tend to have darker-colored points, corresponding to higher AAs, with BISON leading the way, followed by LODE, PCR and ER-ACE.
Table 2. Ablation study on Split Mini-ImageNet ($\mathcal{M}$=1k). “w/o all” represents vanilla ER with the NCM classifier.
Method DC $\mathcal{L}_{PAL}$ $\mathcal{L}_{Align}$ AA $\uparrow$ AF $\downarrow$ AI $\downarrow$
BISON (ours) $\bm{24.2\pm 0.4}$ $10.0\pm 0.6$ $\bm{5.9\pm 8.7}$
w/o DC & PAF $18.1\pm 1.3$ $\bm{7.4\pm 0.4}$ $14.8\pm 8.4$
w/o PAF & PAL $22.6\pm 0.8$ $11.0\pm 0.7$ $6.5\pm 7.5$
w/o PAF $23.4\pm 0.6$ $10.1\pm 0.8$ $6.5\pm 9.1$
w/o PAL $22.5\pm 0.6$ $10.4\pm 1.0$ $7.3\pm 8.2$
w/o all $18.9\pm 1.1$ $8.1\pm 0.9$ $13.1\pm 8.4$
5.3. Ablation Studies
Impact of Each Component. Table 2 presents the results of our ablation study on Split Mini-ImageNet with 1k buffer samples. To assess the impact of different components in the BISON method, we examine performance changes of every possible combinations [^1] Note that PAF cannot solely be used without DC, and hence either ‘w/o DC’ or ‘w/o DC & PAL’ is not a feasible case. of individual components: Dual Classifiers (DC), redesigned Proxy-Anchor Loss (PAL), and Proxy Alignment Feedback (PAF), while commonly using the NCM classifier for inference. We first verify the impact of DC by removing DC and PAF together, as PAF is a backward transfer strategy that depends on DC and cannot function independently. The results, indicated by the reduced AA and increased AI in the ‘w/o DC & PAF’ setting, demonstrate the critical role of DC in enhancing plasticity as well as accuracy. This is partly because removing DC also includes discarding its underlying structure using cosine normalization. The comparison between BISON and ‘w/o PAF’ shows that PAF further improves both plasticity and accuracy when combined with DC. Removing PAL also results in a decline in both stability and plasticity corresponding to increased AF and AI values in ‘w/o PAL’, leading to a decrease in AA. The ‘w/o all’ setting, which represents vanilla ER with NCM, exhibits the worst overall performance. This indicates that every component is essentially required to obtain the final performance and to maintain a good balance between AF and AI.